# Multiscale scattered data analysis in samplet coordinates

### v1

@@ 7. Conclusion
We have proposed an efficient multiscale approach for scattered data interpolation with globally supported RBFs in combination with the samplet matrix compression. In particular, we have extended existing results on the boundedness of the condition numbers of the generalized Vandermonde matrices within the multiscale approximation towards globally supported RBFs. For appropriately chosen lengthscale parameters, the condition numbers stay independent of the level, which allows for an efficient solution of the resulting linear system by the conjugate gradient method with a bounded number of iterations for a given accuracy. The resulting algorithm has an overall cost of $\mathcal{O}(N\log N)$ for $N$ quasi-uniform data sites, given that the points per level form a geometric sequence. The numerical results demonstrate that the presented approach achieves a similar accuracy to the multiscale approach by compactly supported RBFs, while overcoming previous limitations that made the use of globally supported RBFs impractical for large-scale problems. We have demonstrated that the present multiscale approach can handle datasets with millions of data in both two and three spatial dimensions. As test cases, we have considered smooth and non-smooth data in two dimensions and a complex geometry in three spatial dimensions. In all cases, we achieved accurate results that would have been computationally infeasible without the samplet matrix compression.
Acknowledgement. SA and MM were funded by the SNSF starting grant “Multiresolution methods for unstructured data” (TMSGI2_211684).

### v2

@@ 8. Conclusion
We have proposed an efficient multiscale approach for scattered data interpolation with globally supported RBFs in combination with the samplet matrix compression. In particular, we have extended existing results on the boundedness of the condition numbers of the generalized Vandermonde matrices within the multiscale approximation towards globally supported RBFs. For appropriately chosen lengthscale parameters, the condition numbers of the diagonal blocks of the multiscale system stay bounded independently of the level. As a consequence, the diagonally scaled multiscale system becomes well conditioned. We have exploit this fact and derived a general error estimate bounding the consistency error issuing from a numerical approximation of the multiscale system. This result particularly allows to rigorously bound the consistency error arising from the samplet matrix compression. The sparse representation of the multiscale system in samplet coordinates allows for an efficient solution of the linear system by the conjugate gradient method with a bounded number of iterations for a given accuracy, due to the well conditioning. The resulting algorithm has an overall cost of $\mathcal{O}(N\log^{2}N)$ for $N$ quasi-uniform data sites, given local approximation spaces with exponentially decreasing dimension with decreasing level. The numerical results demonstrate that the presented approach achieves a similar accuracy to the multiscale approach using compactly supported RBFs, as presented in [29], where the corresponding experiments can be found. Meanwhile, our approach overcomes previous limitations that made the use of globally supported RBFs impractical for large-scale problems. We have demonstrated that the present multiscale approach can handle datasets with millions of data in both two and three spatial dimensions. As test cases, we have considered smooth and non-smooth data in two dimensions and a complex geometry in three spatial dimensions. In all cases, we achieved accurate results that would have been computationally infeasible without the samplet matrix compression.